# User Simulation in the Era of Generative AI: User Modeling, Synthetic Data Generation, and System Evaluation

### v1

@@ 7. User Simulation as a Step toward AGI
The overarching goal of developing a realistic user simulator is, in many respects, aligned with the broader objective of creating intelligent agents with human-like intelligence, namely, the pursuit of AGI. A sophisticated user simulator that mimics human behaviour would respond to system interactions and select appropriate actions to accomplish a given task. In a similar way, an intelligent agent designed to assist a user in completing a task would respond to the user’s actions and select appropriate actions (such as choosing suitable system services to support the user). Thus, both user simulation agents and intelligent task agents share a common modeling approach, enabling optimization through reinforcement learning algorithms. The key differences lie in the definitions of states, actions, and reward functions. For user simulation, the goal (reward) is to mimic real user behaviour, while for task agents, the reward is typically tied to providing an effective and user-friendly service for task completion. Additionally, user simulation centers on user actions, whereas intelligent agents focus on system responses. Because of this similarity, the technical challenges in building intelligent user simulators may mirror those encountered in developing intelligent task agents. Consequently, we anticipate broader connections between research in user simulation and various other fields of AI.
The Importance of User Simulation in Human-AI Collaboration
Realizing the full potential of human-AI collaboration requires more than just superhuman performance from AI agents. To truly maximize combined intelligence, these agents must also account for the inherent variability in human behaviour, including suboptimal actions, diverse problem-solving approaches, and individual preferences. Recent work in chess, a domain that has long served as a Petri dish for AI research, underscores this challenge. Studies have shown that human players paired with AI agents tailored to their skill level outperform those partnered with more powerful AI agents that are not adjusted for skill-compatibility (Hamade et al., 2024). This suggests that simply providing the most powerful AI assistance may not be optimal. Instead, simulating the user’s knowledge, preferences, and decision-making processes, and adjusting the AI’s assistance accordingly can lead to more effective collaboration. [^1] Note that due to legal responsibility, even autonomous AI systems must collaborate with human owners to minimize risk and ensure safety. This requires the AI to simulate the human owner’s value system and trade-off preferences. Such an approach requires user simulation agents to be closely integrated with task agents, allowing the task agent to leverage the user simulation agent to receive feedback and optimize its interaction policy. Likewise, the user simulation agent must adapt to model how new AI systems behave over time. Therefore, the interdependence between research on intelligent task agents and user simulation is inherent and may persist until AGI is achieved.
Large Language Models as Building Blocks for Intelligent Agents
The emergence of LLMs could accelerate this integration and synergy, as these models may very well serve as a foundational building block for both types of agents. Indeed, LLMs have already fueled their extensive adoption as both task agents and simulation tools in various domains and applications. For comprehensive surveys of work combining agent-based modeling and LLMs across various fields (social science, natural science, and engineering) and domains (e.g., cyber, physical, social environment) we refer readers to (Gao et al., 2024) and (Wang et al., 2024a).
As the uses of LLMs for building user simulation agents and task agents are rapidly expanding, it is also crucial to recognize their limitations. LLM-generated responses can be unpredictable and sometimes unsafe, causing concerns about the safety of an intelligent task agent when deployed in the real world and the realisticity of a user simulation agent. For user simulation, current LLMs would inevitably exhibit unrealistic or incoherent behaviours and may also lack the natural variation observed in real human interactions. Furthermore, LLMs often possess more knowledge than average humans and generate overly “perfect” responses, which may not be a problem for an intelligent task agent (indeed, it could be a feature!), but would likely lead to the simulation of unrealistic “superusers” by a user simulation agent. While prompting techniques can guide LLM behaviour, ensuring strict adherence to instructions remains a challenge (Balog and Zhai, 2024b). Importantly, more fundamental challenges suggest that LLMs alone will not be sufficient. Critically, LLMs have insufficient knowledge of human behaviour. While they might be aware of concepts like patience or satisfaction, they lack the training data to model the human dynamics of such behaviours. Similarly, LLMs lack a deep understanding of human cognitive processes, such as decision-making, memory recall, and attention span. They may fail to accurately simulate phenomena like cognitive biases or the limitations of working memory. These shortcomings hinder their ability to generate realistic simulations of human users. To overcome this limitation, LLMs must be extended with components that capture a wider range of human cognitive abilities. The human brain is known to consist of two distinct systems: System 1, which is intuitive, fast, but not reliable, and System 2, which is logical, deliberate, but slower (Daniel, 2013). While current LLMs appear to be able to simulate System 1 well, they lack the deliberate reasoning and planning capabilities characteristic of System 2. Integrating the functionality of System 2 into LLMs is crucial for more realistic user simulation. In this direction, neurosymbolic approaches, which combine neural networks with symbolic reasoning, represent a promising future direction (Garcez and Lamb, 2023).
@@ 8. Conclusion and Outlook
User simulation plays an increasingly important role in the era of Generative AI with a wide range of applications, including user behaviour modeling and analysis, data augmentation, and system evaluation. Each application area may require different simulation techniques and will likely lead to the development of specialized simulators. At the same time, there is clear synergy among the different lines of exploration and the various research directions can enhance and inform one another. Over time, these diverse lines of research may even converge as we advance toward AGI. Much of the research today on user simulation is scattered across different communities. We envision a convergence of research in this area in the future, where researchers from multiple communities and disciplines would have increasingly more engagement and collaboration. The user simulation portal at http://usersim.ai is an emerging platform with the potential to facilitate this, though further effort is required to accelerate interdisciplinary research and foster cross-disciplinary collaboration. In the short term, progress can be made by regularly hosting workshops on user simulation at different major conferences of different communities, which would facilitate the creation and growth of an interdisciplinary research community around the general topic of user simulation.
To facilitate validation of user simulators at scale, it is essential to involve industry partners who have access to real user behaviour data. Since industry partners would benefit from access to user simulators relevant to their businesses, there may be incentive for them to collaborate with academic researchers on establishing a sustainable ecosystem for accelerating research, development, and applications of user simulation. For example, user simulators built by academic researchers can be made available via open source and can then be validated by industry partners using their real user data. The validation results can be made publicly available to facilitate assessment of progress in developing effective user simulators, which in turn would facilitate development of increasingly more effective user simulators, enabling many applications (including modeling and analyzing real-world user behaviour, evaluation of commercial application systems using user simulation, and data augmentation for training AI systems), forming a positively reinforced and potentially self-sustainable innovation ecosystem, as shown in Figure 3.

### v2

@@ 6.5. Software Systems
7. Ethical Considerations: Bias, Fairness, and Safety
As user simulation becomes increasingly integral to modeling, data augmentation, and system evaluation, it is critical to proactively address the ethical implications inherent in mimicking human behaviour. If left unchecked, biases embedded within a user simulator will inevitably propagate into the synthetic data it generates, creating a feedback loop that degrades the safety and fairness of any downstream AI models trained on that data. However, while user simulation introduces potential ethical risks, it concurrently offers a powerful mechanism to address these exact challenges in broader AI development. Rather than merely reflecting existing inequities, controlled simulation serves as a proactive tool to reduce bias, guarantee fair representation, and systematically stress-test system safety. Table 4 summarizes these core ethical dimensions, contrasting the inherent risks with the proactive opportunities simulation affords.
Table 4. Summary of the ethical dimensions in user simulation, contrasting inherent risks with proactive mitigation opportunities.
Ethical Inherent risk Proactive opportunity
Dimension (if unmitigated) (simulation as a tool)
Bias & data skew Simulators inherit and amplify societal biases or interaction skews present in uncorrected training corpora or historical logs Enables explicit data debiasing; simulators can generate balanced synthetic datasets to correct historical gaps before downstream training
Fairness & representation Relying on a monolithic, “average” simulated user marginalizes minority groups and obscures disparate system performance Facilitates disaggregated evaluation; persona-based simulation ensures rigorous testing against underrepresented groups
Safety & guardrails Simulators lacking proper constraints may inadvertently generate toxic, harmful, or unsafe behaviours during system interaction Unlocks scalable red-teaming; adversarial user simulation intentionally stress-tests system vulnerabilities and safety policies without risking human harm
7.1. Bias and Data Debiasing
Generative foundation models are trained on vast, unfiltered internet corpora, meaning they inherently capture and reproduce historical societal biases (Bender et al., 2021) and often disproportionately reflect the opinions, values, and ideologies of specific dominant demographic groups rather than a globally representative population (Santurkar et al., 2023). Similarly, data-driven simulators inherit the interaction biases present in the logged historical data they learn from. When these biased models are used to power a user simulator, there is a significant risk that the simulated interactions will skew toward stereotypes, over-represent dominant demographics, or provide skewed feedback to the system being evaluated. However, precisely because a simulator is a controllable agent, it can be utilized as a tool for bias mitigation. For example, early approaches to debiasing simulators demonstrated that correcting logged historical data prior to simulation effectively prevents existing interaction biases from degrading downstream policy optimization (Huang et al., 2020). Furthermore, in scenarios where real-world interaction logs are heavily skewed (e.g., containing primarily data from younger, tech-savvy demographics), a user simulator can be parameterized to artificially generate balanced synthetic datasets. By generating interactions that intentionally compensate for historical data gaps, developers can debias the training pipelines of downstream AI systems, ensuring they learn from a more equitable distribution of user behaviours.
7.2. Fairness and Diverse Representation
Algorithmic fairness requires that an AI system performs equitably across all user demographics, regardless of language, culture, or socioeconomic status. A common pitfall in system evaluation is relying on a monolithic, “average” simulated user, which effectively marginalizes minority groups and obscure failure modes that only affect specific subpopulations. To counteract this, researchers can leverage persona-based user simulation. These approaches span a broad spectrum: ranging from prompting models with specific demographic variables (Argyle et al., 2023) to conditioning them on richly detailed narrative backstories that capture more nuanced, intersectional identities (Moon et al., 2024). By deliberately instantiating an array of simulated users that reflect diverse cultural backgrounds and varying levels of domain expertise, developers can conduct rigorous, disaggregated evaluations. This ensures that the AI system is explicitly tested against underrepresented or vulnerable user groups, allowing developers to identify and rectify performance disparities before real-world deployment.
7.3. Safety and Adversarial Simulation
The deployment of interactive AI systems carries inherent safety risks, ranging from the generation of toxic content to susceptibility to prompt injection attacks. If a user simulator lacks proper guardrails, it may inadvertently generate harmful behaviours during the training phase. Conversely, this capacity for negative behaviour can be harnessed intentionally for “red teaming”—the practice of using AI to safely attack and test other AI systems (Perez et al., 2022). By configuring a user simulator to act as an adversarial user (e.g., one attempting to bypass safety filters, extract private information, or induce system errors), developers can expose edge cases and vulnerabilities at scale. In this context, user simulation becomes an indispensable security tool, allowing researchers to stress-test the boundaries of an AI system’s safety policies in a secure, simulated environment without exposing real users to harm.
@@ 8. User Simulation as a Step toward AGI
The overarching goal of developing a realistic user simulator is, in many respects, fundamentally aligned with the broader objective of creating intelligent agents with human-like intelligence, namely, the pursuit of AGI. A critical bottleneck hindering faster progress toward AGI is the heavy reliance on human interaction data for training and evaluation. This process is inherently slow, expensive, and difficult to scale. User simulation addresses this directly: by generating synthetic interaction data and providing scalable evaluation environments, it acts not merely as a useful tool, but as a critical catalyst required to accelerate AGI development (Balog and Zhai, 2026). Furthermore, this alignment is reflected in shared technological foundations. Importantly, this is not a recent trend, but an empirical reality evident throughout AI history. From early rule-based expert systems to probabilistic models, and now to the current wave of Transformer architectures and LLMs, advancements in core AI technologies have consistently been leveraged simultaneously to build both more capable task agents and more realistic user simulators. Today, this structural alignment is most evident in how both user simulation agents and intelligent task agents share a common mathematical modeling approach, enabling optimization through reinforcement learning algorithms. The key differences lie in the definitions of states, actions, and reward functions. For user simulation, the goal (reward) is to mimic real user behaviour, while for task agents, the reward is typically tied to providing an effective and user-friendly service for task completion. Because of this deep synergy, the technical challenges in building intelligent user simulators mirror those encountered in developing intelligent task agents. Consequently, we anticipate that research into user simulation technology and intelligent task agents must advance hand-in-hand. The necessity of this co-evolution becomes particularly evident when examining the architectural requirements for true human-AI collaboration.
Figure 3. Conceptual “Theory of Mind” architecture for an advanced interactive AI system. The task agent utilizes an embedded user simulation agent as an internal world model. This creates a cognitive feedback loop where the decision policy proposes candidate actions, the user simulator forecasts the human’s likely reaction, and the policy optimizes its choice before executing a final action in the real environment.
8.1. The Importance of User Simulation in Human-AI Collaboration
Realizing the full potential of human-AI collaboration requires more than just superhuman performance from AI agents. To truly maximize combined intelligence, these agents must also account for the inherent variability in human behaviour, including suboptimal actions, diverse problem-solving approaches, and individual preferences. Recent work in chess, a domain that has long served as a Petri dish for AI research, underscores this challenge. Studies have shown that human players paired with AI agents tailored to their skill level outperform those partnered with more powerful AI agents that are not adjusted for skill-compatibility (Hamade et al., 2024). This suggests that simply providing the most powerful AI assistance may not be optimal. Instead, simulating the user’s knowledge, preferences, and decision-making processes, and adjusting the AI’s assistance accordingly can lead to more effective collaboration. [^1] Note that due to legal responsibility, even autonomous AI systems must collaborate with human owners to minimize risk and ensure safety. This requires the AI to simulate the human owner’s value system and trade-off preferences. Such an approach requires user simulation agents to be closely integrated with task agents, allowing the task agent to leverage the user simulation agent to receive feedback and optimize its interaction policy. In essence, an autonomous intelligent agent that leverages its own internal user simulation agent to anticipate human needs and optimize interactions is implementing a functional analogue of Theory of Mind (Premack and Woodruff, 1978). Figure 3 illustrates this conceptual architecture. Rather than relying on purely reactive heuristics, an advanced task agent embeds a user simulation agent directly within its cognitive framework to act as an internal world model (LeCun, 2022; Ding et al., 2025). During interaction, the task agent’s decision policy does not immediately output a response to the real human user. Instead, it initiates an internal feedback loop: the policy proposes a candidate action, and the user simulator forecasts the human’s likely response. By mentally “rehearsing” these interactions through simulated trial-and-error, the system can optimize its planning phase, ultimately executing actions in the real environment that are significantly more aligned, empathetic, and effective. Likewise, the user simulation agent must adapt to model how new AI systems behave over time. Therefore, the interdependence between research on intelligent task agents and user simulation is inherent and may persist until AGI is achieved. For a more comprehensive discussion on the indispensable role of simulation in AGI, we refer the reader to (Balog and Zhai, 2026).
8.2. Large Language Models as Building Blocks for Intelligent Agents
The emergence of LLMs could accelerate this integration and synergy, as these models may very well serve as a foundational building block for both types of agents. Indeed, LLMs have already fueled their extensive adoption as both task agents and simulation tools in various domains and applications. For comprehensive surveys of work combining agent-based modeling and LLMs across various fields (social science, natural science, and engineering) and domains (e.g., cyber, physical, social environment) we refer readers to (Gao et al., 2024) and (Wang et al., 2024b).
As the uses of LLMs for building user simulation agents and task agents are rapidly expanding, it is also crucial to recognize their limitations. LLM-generated responses can be unpredictable and sometimes unsafe, causing concerns about the safety of an intelligent task agent when deployed in the real world and the realisticity of a user simulation agent. For user simulation, current LLMs would inevitably exhibit unrealistic or incoherent behaviours and may also lack the natural variation observed in real human interactions. Furthermore, LLMs often possess more knowledge than average humans and generate overly “perfect” responses, which may not be a problem for an intelligent task agent (indeed, it could be a feature!), but would likely lead to the simulation of unrealistic “superusers” by a user simulation agent. While prompting techniques can guide LLM behaviour, ensuring strict adherence to instructions remains a challenge (Balog and Zhai, 2024b). Importantly, more fundamental challenges suggest that LLMs alone will not be sufficient. Critically, LLMs have insufficient knowledge of human behaviour. While they might be aware of concepts like patience or satisfaction, they lack the training data to model the human dynamics of such behaviours. Similarly, LLMs lack a deep understanding of human cognitive processes, such as decision-making, memory recall, and attention span. They may fail to accurately simulate phenomena like cognitive biases or the limitations of working memory. These shortcomings hinder their ability to generate realistic simulations of human users.
To overcome this limitation, LLMs must be extended with components that capture a wider range of human cognitive abilities. The human brain is known to consist of two distinct systems: System 1, which is intuitive, fast, but not always reliable, and System 2, which is logical, deliberate, and slower (Daniel, 2013). Current LLM-powered agents can simulate the rapid intuition of System 1 relatively well, but simulating System 2—and achieving seamless integration between the two—remains a formidable open challenge. For example, capturing uniquely human behaviours that require deep, deliberate reasoning, such as our capacity to not only use existing tools but invent new ones to solve unprecedented problems, remains a significant hurdle for current architectures (Xu et al., 2025).
@@ 8.2. Large Language Models as Building Blocks for Intelligent Agents
Recent research has attempted to address these gaps through various architectural enhancements for LLM-powered systems. For instance, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) and GraphRAG (Edge et al., 2025) are increasingly used to ground an agent’s responses in external knowledge, with specific retrieval tasks proving important for supporting human-like AI agents (Zhai, 2025). Furthermore, advanced prompting paradigms and agentic workflows—such as Chain-of-Thought (CoT) (Wei et al., 2022), Tree-of-Thoughts (ToT) (Yao et al., 2023b), or the Reflection Pattern (Madaan et al., 2023)—can be employed to elicit multi-step reasoning and dynamically correct LLM errors. Broader integrations of LLMs with cognitive architectures have also been explored to enable models to solve complex, open problems (Sumers et al., 2024).
However, it is crucial to recognize the fundamental limitations of these techniques in the context of user simulation. While RAG, CoT/ToT, and agentic reflection make LLMs better problem solvers, they do not necessarily make them more cognitively plausible. For instance, while a “Critic” agent is highly effective for optimizing a task agent’s performance, applying such a cycle of self-criticism to user simulation is only valid if it is explicitly mapped to a human cognitive process, such as metacognition or conscious self-correction. Human reasoning is heavily constrained by bounded rationality, working memory limits, and cognitive biases. Therefore, simply prompting an LLM to “think harder” or “critique itself” to achieve optimal results does not adequately mimic the reality of human cognitive processes. Instead, achieving true cognitive plausibility requires the seamless integration of System 2 functionality into LLM architectures. Key challenges include determining how symbolic reasoning can be used to explicitly regulate agent behavior (e.g., via reward supervision), and how agents can acquire symbolic knowledge dynamically through interactions with their environment. In this direction, neurosymbolic approaches, which combine neural networks with symbolic reasoning, represent a much more promising future direction to capture the actual mechanisms of human thought (Garcez and Lamb, 2023).
@@ 9. Conclusion and Outlook
User simulation plays an increasingly important role in the era of Generative AI with a wide range of applications, including user behaviour modeling and analysis, data augmentation, and system evaluation. Each application area may require different simulation techniques and will likely lead to the development of specialized simulators. At the same time, there is clear synergy among the different lines of exploration and the various research directions can enhance and inform one another. Over time, these diverse lines of research may even converge as we advance toward AGI. As this interdisciplinary community matures, a critical next step will be the development of unified metamodels, shared architectures, and exchange standards for user simulation. Currently, the lack of such standards hinders interoperability between simulators developed in different subfields. Establishing these frameworks will be essential for addressing the highly complex, multi-angle requirements of user simulation at scale. Much of the research today on user simulation is scattered across different communities. We envision a convergence of research in this area in the future, where researchers from multiple communities and disciplines would have increasingly more engagement and collaboration. The user simulation portal at http://usersim.ai is an emerging platform with the potential to facilitate this, though further effort is required to accelerate interdisciplinary research and foster cross-disciplinary collaboration. In the short term, progress can be made by regularly hosting workshops on user simulation at different major conferences of different communities, which would facilitate the creation and growth of an interdisciplinary research community around the general topic of user simulation.
9.1. Establishing a Sustainable Innovation Ecosystem
To facilitate validation of user simulators at scale, it is essential to involve industry partners who have access to real user behaviour data. Such collaboration can form the basis of a positively reinforced innovation ecosystem, as illustrated in Figure 4. Rather than a one-sided transfer of technology, this ecosystem is designed so that all stakeholders act simultaneously as beneficiaries and contributors. While academic researchers drive innovation by developing and open-sourcing novel user simulators, we acknowledge that expecting industry partners to validate these models without clear economic incentives is idealistic. Validating academic models against proprietary data incurs significant costs, including compute resources, engineering overhead, and strict data privacy compliance. To offset these real-world constraints and ensure the ecosystem is self-sustaining, there must be a clear economic value exchange. With the proposed ecosystem, industry partners may be incentivized to bear these validation costs because the collaboration provides them with vital pre-competitive assets: independent benchmarks to prove their AI systems are safe, standardized evaluation frameworks they do not have to build from scratch, and direct engagement with an academic talent pipeline. Once validated against real-world data, the best simulators become directly useful for evaluating the industry’s own product systems. Meanwhile, this industry-provided validation service would attract academic researchers to the platform to test new ideas and algorithms. This continuous influx of academic innovation improves the overall quality of user simulators, which in turn reinforces the value proposition for industry partners, closing the loop on a self-sustaining collaboration platform.
As a practical first step toward this broader vision, the community can establish a scaled-down “minimum viable ecosystem.” Rather than beginning with complex, large-scale infrastructural commitments, initial progress can be driven through industry-sponsored shared tasks and public data challenges at major AI conferences. A prominent real-world example of this momentum is the newly launched User Simulation track at the Text REtrieval Conference (TREC), [^2] https://trec.usersim.ai organized by the National Institute of Standards and Technology (NIST). In models like this, an organizing body or industry partner releases a carefully anonymized, limited-scope interaction dataset, and academic teams compete to build the most accurate or cognitively plausible simulator. This approach requires significantly lower upfront investment, navigates legal and data privacy concerns more easily, and builds the foundational trust and baseline metrics necessary to validate the ecosystem concept.
Eventually, to scale beyond limited datasets and fully mitigate data privacy constraints on a continuous basis, this validation can transition toward federated evaluation paradigms. Here, the open-source simulator is brought to the company’s secure data environment, and only aggregated, anonymized validation metrics are made publicly available. Supported financially by joint industry consortiums or government grants, these shared validation results facilitate the assessment of progress, creating a positively reinforced and economically viable innovation loop between academia and industry.